# Learning to Insert for Constructive Neural Vehicle Routing Solver

### v2

@@ Abstract
Neural Combinatorial Optimisation (NCO) is a promising learning-based approach for solving Vehicle Routing Problems (VRPs) without extensive manual design. While existing constructive NCO methods typically follow an appending-based paradigm that sequentially adds unvisited nodes to partial solutions, this rigid approach often leads to suboptimal results. To overcome this limitation, we explore the idea of insertion-based paradigm and propose Learning to Construct with Insertion-based Paradigm (L2C-Insert), a novel learning-based method for constructive NCO. Unlike traditional approaches, L2C-Insert builds solutions by strategically inserting unvisited nodes at any valid position in the current partial solution, which can significantly enhance the flexibility and solution quality. The proposed framework introduces three key components: a novel model architecture for precise insertion position prediction, an efficient training scheme for model optimization, and an advanced inference technique that fully exploits the insertion paradigm’s flexibility. Extensive experiments on both synthetic and real-world instances of the Travelling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP) demonstrate that L2C-Insert consistently achieves superior performance across various problem sizes.
@@ 1 Introduction
Among NCO methods, constructive approaches have attracted considerable research interest [6, 9, 19, 10, 13, 14, 20, 21, 22]. These methods typically employ an appending-based constructive paradigm that sequentially selects unvisited nodes and appends them to the end of the current partial solution. This paradigm is inspired by the sequence-to-sequence framework [23] in natural language processing (NLP), where the model encodes an input sequence into an embedding and then autoregressively generates an output sequence based on this representation. Vinyals et al. [6] pioneered this adaptation through their Pointer Networks (Ptr-Nets), which established the first sequence-to-sequence framework specifically for neural combinatorial optimization. Subsequent research has further advanced constructive NCO methods within this appending-based paradigm, with improvements focusing on model training techniques [7] and architectural enhancements [9]. However, appending-based methods suffer from a fundamental limitation due to their strict operational constraint that only permits adding new nodes to the end of the current partial solution. This constraint prevents potentially more effective modifications and often leads to suboptimal partial solutions caused by early greedy decisions.
@@ 3.1 Model Structure
$\mathbf{e}_{\pi_{i}}$ $=W_{2}\begin{cases}[\mathbf{h}_{\pi_{i}},\mathbf{h}_{\pi_{i+1}}]&i\in\{1:t-2\},\\ [\mathbf{h}_{\pi_{t-1}},\mathbf{h}_{\pi_{1}}],&i=t-1,\end{cases}$ (1)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents the horizontal concatenation operator. This operator produces an initial $2d$-dimensional initial insertion position embedding, which is subsequently projected to a $d$-dimensional vector via the trainable matrix $W_{2}\in\mathbb{R}^{2d\times d}$ to maintain dimensional consistency with node embeddings. The set of position embeddings is denoted as $H_{e}=\{\mathbf{e}_{\pi_{i}}\},i\in\{1:t-1\}$.
Then, these embeddings together are refined by $L$ stacked attention layers. The first layer receives input $X^{(0)}=\{\mathbf{h}_{c},H_{e},H_{u}\}$, with each subsequent $j$-th attention layer processing the output from the $(j-1)$-th layer, culminating in the final output $X^{(L)}$. Then, a linear projection and softmax function are applied to $X^{(L)}$ to compute the insertion probabilities of the current node into all valid positions. The masking operation assigns $\infty$- ∞ to invalid positions (those corresponding to the current node and non-visited nodes) to ensure proper probability normalization over valid insertion locations. The inserting probability $p_{i}$ is calculated as:
@@
$a_{i}$ =$=\begin{cases}x_{i}&\forall i\in\{\pi_{1:t-1}\},\\ -\infty&\text{otherwise,}\end{cases}$
@@ 3.3 Insertion-based Local Reconstruction
To overcome this limitation, we propose an insertion-based local reconstruction method that exploits the inherent flexibility of the insertion paradigm through a distance-based deconstruction strategy, which can enable more effective reconstruction and consequently enhance the model performance. The iterative local reconstruction process alternates between destruction and reconstruction phases. (1) Destruction Phase: The solution is destroyed based on the distance-based proximity of nodes. The process begins by randomly selecting a node from the current solution, followed by identification and removal of its $\alpha$-nearest neighbors, which transition to unvisited nodes. The remaining solution segments automatically form a new reduced loop solution. Here, $\alpha$ represents the destruction size. For example, given a TSP10 solution $(\pi_{1},\pi_{2},\ldots,\pi_{10})$ where $\pi_{5}$ is randomly selected with its 3-nearest neighbors $\{\pi_{2},\pi_{7},\pi_{8}\}$, the removal of $\pi_{5}$ and $\{\pi_{2},\pi_{7},\pi_{8}\}$ yields the looped solution $(\pi_{1},\pi_{3},\pi_{4},\pi_{6},\pi_{9},\pi_{10})$. (2) Reconstruction Phase: The model sequentially inserts the unvisited nodes into this partial solution. Once all nodes are re-inserted, the newly generated solution is compared against its predecessor, and the superior solution is retained for subsequent iterations. This iterative process continues until a predefined computational budget (e.g., a maximum number of iterations) is exhausted.
Table 1: Results on synthetic TSP and CVRP instances. Results marked with an asterisk (*) are from the original paper. The best results are bold. OOM: The method triggered an out-of-memory condition.
@@
L2C-Insert ($I$=100) 7.764 0.0038% 47.9m 23.414 1.275% 4.79m 75.603 5.329% 2.58m
L2C-Insert ($I$=200) 7.76336 0.0014% 1.62h 23.339 0.952% 9.52m 74.733 4.117% 4.14m
L2C-Insert ($I$=500) 7.76328 0.0005% 4.0h 23.265 0.633% 22.94m 73.812 2.834% 8.73m
L2C-Insert ($I$=1000) 7.76326 0.0002% 8.0h 23.230 0.481% 46.4m 73.270 2.079% 16.64m
@@
L2C-Insert ($I$=100) 15.700 0.877% 1.0h 43.868 6.574% 12.08m 253.462 12.063% 5.75m
L2C-Insert ($I$=200) 15.667 0.664% 2.0h 43.677 6.113% 24.66m 252.021 11.426% 9.02m
L2C-Insert ($I$=500) 15.640 0.493% 5.3h 43.416 5.478% 1.05h 248.987 10.085% 19.03m
L2C-Insert ($I$=1000) 15.627 0.413% 10.6h 43.152 4.836% 2.05h 245.732 8.646% 34.72m
@@ 4 Experiments
[…]
@@
The embedding dimension is set to $d=128$, the hidden dimension of the feed-forward layer is set to 512, the query dimension $d_{q}$ is set to 16, and the head number in multi-head attention is set to 8. We set the number of attention layers in the decoder to $L=9$ following [14]. Both the TSP and CVRP models are trained on one million instances of size 100. For training, we employ the Adam optimizer [31] with an initial learning rate of 1×10-${}^{\mbox{-}4}$, decaying by a factor of 0.97 after each epoch. The TSP and CVRP models undergo 50 training epochs with a batch size of 1024, respectively. All experiments, including training, testing, and evaluation, are conducted on a single NVIDIA GeForce RTX 4090 GPU with 24 GB of memory to ensure consistent computational conditions.
@@
We compare L2C-Insert with a comprehensive set of baseline methods including (1) Classical Solvers: Concorde [35], LKH3 [42], and HGS [36]; (2) Constructive NCO Methods: POMO [10]; BQ [14], LEHD [13], INViT [43], and ELG [15]. (3) Heatmap-based NCO Methods: Att-GCN+MCTS [37], DIFUSCO [44], and T2T [45]; (4) Two-stage NCO Methods: H-TSP [46], SO-mixed [47], GLOP [34], and UDC [34]; (5) Improvement-based NCO Methods: NeurOpt [48]. We refer to Appendix C for implementation details of these baselines.
@@
For inference, the L2C-Insert model first generates an initial solution through greedy rollout, then progressively refines solution quality via insertion-based local reconstruction. The number of iterations is denoted as “$I$=”. Following [14], we accelerate inference by processing only the $k$-nearest unvisited nodes and positions of the current node, with $k$ set to 100 for TSP and 200 for CVRP. To further accelerate this process, we limit the local reconstruction destruction size $\alpha$ to 300 across all instances. Following [34], to enhance the consistency of inputs for our model on large-scale instances with more than 1K nodes, we employ min-max feature scaling, which adjusts the x-axis and y-axis coordinate values to the interval [0, 1]. A hyperparameter study and discussion on these settings are provided in Appendix D. For performance evaluation, we measure: 1) average tour length (Length), 2) performance gap relative to baseline solvers (Gap), and 3) total inference time (Time). Note that reported inference times for classical CPU-based solvers are not directly comparable to GPU-executed learning-based methods.
@@ 4.1 Comparative Results
Tables 1 and 2 present the comparative results on uniformly distributed TSP and CVRP instances scaling from 100 to 100K nodes. These results show that L2C-Insert consistently achieves outstanding performance across these scales with a reasonable runtime. On TSP100, L2C-Insert ($I$=200) reduces the gap of the advanced NCO baseline LEHD+RRC1000 by 12.5% with a speedup of 1.4×1.4$\times$1.4 ×. On TSP1K, L2C-Insert ($I$=1000) achieves a significant reduction of 34.2% while with a speedup of 6.7$\times$ compared to LEHD+RRC1000. On TSP10K, L2C-Insert ($I$=1000) outperforms the strong heatmap-based solver DIFUSCO and two-stage solver GLOP with 19.4% and 54.3% gap reduction, while with 2.8$\times$ and 1.5$\times$ speedup, respectively. On TSP50K and TSP100K, L2C-Insert ($I$=200) outperforms the two-stage NCO method GLOP with a gap reduction of 24.8% and 4.3%, respectively. On CVRP100, L2C-Insert ($I$=1000) outperforms LEHD RRC1000 by 1.7%, and on CVRP1K it achieves a competitive gap while with a speedup of 8.7$\times$. On CVRP10K, L2C-Insert ($I$=1000) outperforms LEHD RRC1000 with a speed-up of 14$\times$. On CVRP50K and CVRP100K, L2C-Insert using greedy rollout achieves significant gap reductions of 62.7% and 71.4% compared to INViT, respectively. Overall, our method shows good scalability and can achieve outstanding performance across small-scale to very large-scale problem instances with up to 100K nodes. The experimental results on real-world TSPLIB and CVRPLIB instances are provided in Table 3. These results show that L2C-Insert consistently outperforms representative neural solvers on all instance groups, demonstrating robust generalization ability.
@@ 4.2 Ablation Study
We conduct a comparative evaluation between two constructive NCO approaches: learning to append (L2C-Append) and learning to insert (L2C-Insert). While L2C-Append shares a similar model architecture with L2C-Insert, it utilizes an appending construction framework akin to LEHD [13]. Both models are trained on TSP100 under identical budgets and subsequently evaluated on TSP100, TSP1K, and TSP10K datasets. For inference, we apply local reconstruction with 200 and 1000 iterations using the aforementioned settings. As detailed in Table 4, L2C-Insert, with only 200 reconstruction iterations, outperforms L2C-Append even when the latter uses 1000 iterations of local reconstruction while also consuming less runtime. Remarkably, L2C-Insert ($I$=1000) outperforms L2C-Append ($I$=1000) by achieving substantial gap reductions of at least 50.6% on these datasets. These findings highlight that the L2C-Insert framework’s greater flexibility in solution construction directly contributes to its superior performance.
@@
We ablate the effect of distance-based strategies on local reconstruction. Specifically, on TSP1K and TSP10K, we test the L2C-Insert model using local reconstruction ($I$=1000) with the distance-based and sequence-based destruction separately. As demonstrated in Table 5, the distance-based destruction strategy achieves substantially better performance, indicating its greater effectiveness in helping local reconstruction escape local optima.
@@ Conclusion
In this work, we propose L2C-Insert, a novel learning to insert framework to explore the potential of the insertion paradigm for constructive NCO. By strategically inserting unvisited nodes at any valid position within the partial solution, L2C-Insert can flexibly construct high-quality solutions for routing problems. Extensive comparisons with advanced neural solvers demonstrate L2C-Insert’s outstanding performance on TSP and CVRP.
@@
Although L2C-Insert has demonstrated outstanding performance, its performance could be further boosted through the development of more effective model architectures. These improved architectures would better capture the intricate relationships between current nodes, unvisited nodes, and positions within partial solutions. Extending the L2C-Insert framework to address a broader range of combinatorial optimization problems is also a promising research direction in future work.

### v3

@@ Abstract
Neural Combinatorial Optimisation (NCO) is a promising learning-based approach for solving Vehicle Routing Problems (VRPs) without extensive manual design. While existing constructive NCO methods typically follow an appending-based paradigm that sequentially adds unvisited nodes to partial solutions, this rigid approach often leads to suboptimal results. To overcome this limitation, we explore the idea of the insertion-based paradigm and propose Learning to Construct with Insertion-based Paradigm (L2C-Insert), a novel learning-based method for constructive NCO. Unlike traditional approaches, L2C-Insert builds solutions by strategically inserting unvisited nodes at any valid position in the current partial solution, which can significantly enhance the flexibility and solution quality. The proposed framework introduces three key components: a novel model architecture for precise insertion position prediction, an efficient training scheme for model optimization, and an advanced inference technique that fully exploits the insertion paradigm’s flexibility. Extensive experiments on both synthetic and real-world instances of the Travelling Salesman Problem (TSP) and Capacitated Vehicle Routing Problem (CVRP) demonstrate that L2C-Insert consistently achieves superior performance across various problem sizes. The code is available at https://github.com/CIAM-Group/L2C_Insert.
@@ 1 Introduction
Among NCO methods, constructive approaches attract considerable research interest [6, 19, 20, 21, 22, 23, 24, 25, 26]. These methods typically employ an appending-based constructive paradigm that sequentially selects unvisited nodes and appends them to the end of the current partial solution. This paradigm is inspired by the sequence-to-sequence framework [27] in natural language processing (NLP), where the model encodes an input sequence into an embedding and then autoregressively generates an output sequence based on this representation. Vinyals et al. [6] pioneer this adaptation through their Pointer Networks (Ptr-Nets), which establishes the first sequence-to-sequence framework specifically for neural combinatorial optimization. Subsequent research has further advanced constructive NCO methods within this appending-based paradigm, with improvements focusing on model training techniques [7] and architectural enhancements [19]. However, appending-based methods suffer from a fundamental limitation due to their strict operational constraint that only permits adding new nodes to the end of the current partial solution. This constraint prevents potentially more effective modifications and often leads to suboptimal partial solutions caused by early greedy decisions.
@@ 3.1 Model Structure
[…]
where $\cdot$$\cdot$] represents the horizontal concatenation operator. This operator produces an initial $2d$-dimensional initial insertion position embedding, which is subsequently projected to a $d$-dimensional vector via the trainable matrix $W_{2}\in\mathbb{R}^{2d\times d}$ to maintain dimensional consistency with node embeddings. The set of position embeddings is denoted as $H_{e}=\{\mathbf{e}_{\pi_{i}}\},i\in\{1:t-1\}$.
Then, these embeddings together are refined by $L$ stacked attention layers. The first layer receives input $X^{(0)}=\{\mathbf{h}_{c},H_{e},H_{u}\}$, with each subsequent $j$-th attention layer processing the output from the (j1)(j-1)-th layer, culminating in the final output $X^{(L)}$. Then, a linear projection and softmax function are applied to $X^{(L)}$ to compute the insertion probabilities of the current node into all valid positions. The masking operation assigns -$\infty$ to invalid positions (those corresponding to the current node and non-visited nodes) to ensure proper probability normalization over valid insertion locations. The inserting probability $p_{i}$ is calculated as:
@@
[…]
@@ 3.3 Insertion-based Local Reconstruction
To overcome this limitation, we propose an insertion-based local reconstruction method that exploits the inherent flexibility of the insertion paradigm through a distance-based deconstruction strategy, which can enable more effective reconstruction and consequently enhance the model performance. The iterative local reconstruction process alternates between destruction and reconstruction phases. (1) Destruction Phase: The solution is destroyed based on the distance-based proximity of nodes. The process begins by randomly selecting a node from the current solution, followed by the identification and removal of its $\alpha$-nearest neighbors, which transition to unvisited nodes. The remaining solution segments automatically form a new reduced loop solution. Here, $\alpha$ represents the destruction size. For example, given a TSP10 solution $(\pi_{1},\pi_{2},\ldots,\pi_{10})$ where $\pi_{5}$ is randomly selected with its 3-nearest neighbors $\{\pi_{2},\pi_{7},\pi_{8}\}$, the removal of $\pi_{5}$ and $\{\pi_{2},\pi_{7},\pi_{8}\}$ yields the looped solution $(\pi_{1},\pi_{3},\pi_{4},\pi_{6},\pi_{9},\pi_{10})$. (2) Reconstruction Phase: The model sequentially inserts the unvisited nodes into this partial solution. Once all nodes are re-inserted, the newly generated solution is compared against its predecessor, and the superior solution is retained for subsequent iterations. This iterative process continues until a predefined computational budget (e.g., a maximum number of iterations) is exhausted.
Table 1: Results on synthetic TSP and CVRP instances. Results marked with an asterisk (*) are from the original paper. The best results are bold. OOM: The method triggered an out-of-memory condition. The reported inference time is the total time required to solve the entire test set.
@@
L2C-Insert (II=100) 7.764 0.0038% 47.9m 23.414 1.275% 4.79m 75.603 5.329% 2.58m
L2C-Insert (II=200) 7.76336 0.0014% 1.62h 23.339 0.952% 9.52m 74.733 4.117% 4.14m
L2C-Insert (II=500) 7.76328 0.0005% 4.0h 23.265 0.633% 22.94m 73.812 2.834% 8.73m
L2C-Insert (II=1000) 7.76326 0.0002% 8.0h 23.230 0.481% 46.4m 73.270 2.079% 16.64m
@@
L2C-Insert (II=100) 15.700 0.877% 1.0h 43.868 6.574% 12.08m 253.462 12.063% 5.75m
L2C-Insert (II=200) 15.667 0.664% 2.0h 43.677 6.113% 24.66m 252.021 11.426% 9.02m
L2C-Insert (II=500) 15.640 0.493% 5.3h 43.416 5.478% 1.05h 248.987 10.085% 19.03m
L2C-Insert (II=1000) 15.627 0.413% 10.6h 43.152 4.836% 2.05h 245.732 8.646% 34.72m
@@ 4 Experiments
Following the standard data generation procedure from prior work [19], we generate datasets for both TSP and CVRP problems. (1) Training Datasets: For each problem (TSP and CVRP), we create training datasets consisting of one million 100-node instances. We obtain (near-)optimal solutions for these instances using Concorde [41] for TSP and HGS [42] for CVRP. (2) Test Datasets: Consistent with established practices [43, 22, 44, 45], we generate five synthetic test datasets with instance sizes ranging from 100 to 100K nodes (100, 1K, 10K, 50K, and 100K). We denote TSP and CVRP instances with $N$ nodes as TSPNN and CVRPNN (e.g., TSP100, CVRP100). The dataset composition is as follows: For TSP, the TSP100 and TSP1K datasets contain 10,000 and 128 instances, respectively, while larger-scale datasets (10K, 50K, 100K nodes) each have 16 instances. CVRP datasets follow a similar structure, with the exception that CVRP1K contains 100 instances, as specified in [44]. Following Hou et al. [44], we set vehicle capacities to 50 for CVRP100, 200 for CVRP1K, and 300 for larger-scale instances. To evaluate our method on real-world large-scale instances, we also collect 49 symmetric TSP instances and 100 CVRP instances with EUC_2D features and $\leq$1K nodes from TSPLIB [46] and CVRPLIB Set-X [47], respectively.
@@
The embedding dimension is set to $d=128$, the hidden dimension of the feed-forward layer is set to 512, the query dimension $d_{q}$ is set to 16, and the head number in multi-head attention is set to 8. We set the number of attention layers in the decoder to $L=9$ following [23]. Both the TSP and CVRP models are trained on one million instances of size 100. For TSP model training, we employ the Adam optimizer [37]. For CVRP model training, we employ the AdamW [48] optimizer, as we find it to be more stable. The initial learning rate is 1×10-$^{\mbox{-}4}$, decaying by a factor of 0.97 after each epoch. The TSP and CVRP models undergo 50 and 15 training epochs, respectively, with a batch size of 1024. All experiments, including training, testing, and evaluation, are conducted on a single NVIDIA GeForce RTX 4090 GPU with 24 GB of memory to ensure consistent computational conditions.
@@
We compare L2C-Insert with a comprehensive set of baseline methods including (1) Classical Solvers: Concorde [41], LKH3 [49], and HGS [42]; (2) Constructive NCO Methods: POMO [21]; BQ [23], LEHD [22], INViT [50], and ELG [30]. (3) Heatmap-based NCO Methods: Att-GCN+MCTS [43], DIFUSCO [51], and T2T [52]; (4) Two-stage NCO Methods: H-TSP [53], SO-mixed [54], GLOP [40], and UDC [40]; (5) Improvement-based NCO Methods: NeurOpt [55]. In line with the established practice in the NCO literature [19, 21], the classical solvers are primarily included to provide a strong, near-optimal baseline for solution quality, rather than to serve as direct competitors in terms of runtime. We refer to Appendix C for implementation details of these baselines.
@@
For inference, the L2C-Insert model first generates an initial solution through greedy rollout, then progressively refines solution quality via insertion-based local reconstruction. The number of iterations is denoted as “II=”. Following [23], we accelerate inference by processing only the $k$-nearest unvisited nodes and positions of the current node, with $k$ set to 100 for TSP and 200 for CVRP. To further accelerate this process, we limit the local reconstruction destruction size $\alpha$ to 300 across all instances. Following [40], to enhance the consistency of inputs for our model on large-scale instances with more than 1K nodes, we employ min-max feature scaling, which adjusts the x-axis and y-axis coordinate values to the interval [0, 1]. A hyperparameter study and discussion on these settings are provided in Appendix D. For performance evaluation, we measure: 1) average tour length (Length), 2) performance gap relative to baseline solvers (Gap), and 3) total inference time (Time). Note that reported inference times for classical CPU-based solvers are not directly comparable to GPU-executed learning-based methods.
@@ 4.1 Comparative Results
[…]
@@ 4.2 Ablation Study
We conduct a comparative evaluation between two constructive NCO approaches: learning to append (L2C-Append) and learning to insert (L2C-Insert). While L2C-Append shares a similar model architecture with L2C-Insert, it utilizes an appending construction framework akin to LEHD [22]. Both models are trained on TSP100 under identical budgets and subsequently evaluated on TSP100, TSP1K, and TSP10K datasets. For inference, we apply local reconstruction with 200 and 1000 iterations using the aforementioned settings. As detailed in Table 4, L2C-Insert, with only 200 reconstruction iterations, outperforms L2C-Append even when the latter uses 1000 iterations of local reconstruction while also consuming less runtime. Remarkably, L2C-Insert (II=1000) outperforms L2C-Append (II=1000) by achieving substantial gap reductions of at least 50.6% on these datasets. These findings highlight that the L2C-Insert framework’s greater flexibility in solution construction directly contributes to its superior performance.
@@
We ablate the effect of distance-based strategies on local reconstruction. Specifically, on TSP1K and TSP10K, we test the L2C-Insert model using local reconstruction (II=1000) with the distance-based and sequence-based destruction separately. As demonstrated in Table 5, the distance-based destruction strategy achieves substantially better performance, indicating its greater effectiveness in helping local reconstruction escape local optima.
@@ Conclusion
In this work, we have proposed L2C-Insert, a novel learning to insert framework to explore the potential of the insertion paradigm for constructive NCO. By strategically inserting unvisited nodes at any valid position within the partial solution, L2C-Insert can flexibly construct high-quality solutions for routing problems. Extensive comparisons with advanced neural solvers demonstrate L2C-Insert’s outstanding performance on TSP and CVRP.
@@
First, although L2C-Insert has demonstrated outstanding performance, its performance could be further boosted through the development of more effective model architectures. These improved architectures would better capture the intricate relationships between current nodes, unvisited nodes, and positions within partial solutions. Second, designing a hybrid NCO solver that combines both appending and inserting moves is a highly promising research direction. Such a model has the potential to be more flexible and powerful than a model restricted to a single type of move. In addition, a limitation may be the greedy, non-learned nearest-neighbor heuristic used for node selection. A promising research direction is to develop a learned policy that works synergistically with the insertion model, potentially through joint or two-stage training. Finally, extending the L2C-Insert framework to address a broader range of combinatorial optimization problems is another promising research direction for future work.
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Natural Science Foundation of China (Grant 62476118, Grant 12426305), the Natural Science Foundation of Guangdong Province (Grant 2024A1515011759), the Guangdong Science and Technology Program (Grant 2024B1212010002), the Center for Computational Science and Engineering at Southern University of Science and Technology, the National Key Research and Development Program of China (Grant 2022YFA1004201), and the Research Grants Council of the Hong Kong Special Administrative Region, China (GRF Project No. CityU 11212524).
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations of the work in Section 5.
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [N/A]
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  19. Justification: We provide the detailed experiment setting in Section 4, hyperparameter settings in Appendix D, and implementation in Appendix B, which is fully sufficient for reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  24. Justification: See the abstract for the link to the code and trained models.
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  29. Justification: We provide the detailed experiment setting in Section 4, hyperparameter settings in Appendix D, and implementation in Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [No]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: Yes
  39. Justification: We describe the compute resources in Section 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  49. Justification: As a learning-based method for solving VRPs, the proposed method does not have any specific potential negative societal impact. Detailed discussion can be found in Appendix O. discussion.
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  59. Justification: We have properly cited the assets (e.g., code, data, models) in section P.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines: